A compound train consists of six gears. The number of teeth on the gears are as follows:


Gear : AB C DE F


No. of teeth : 60 40 50 25 30 24


The gears B and C are on one shaft while the gears D and E are on another shaft. The gear A drives gear B, gear C drives gear D and gear E drives gear F. If the gear A transmits 1.5 kW at 100 r.p.m. and the gear

train has an efficiency of 80 per cent, find the torque on gear F.


1
Expert's answer
2021-05-31T06:20:31-0400

Given that :

"T_A=60"

"T_A=40"

"T_C=50"

"T_D=25"

"T_E=30"

"T_F=24"

Velocity ratio (VR)"=\frac{No \space of \space teeth \space on \space driven \space gears }{No \space of \space teeth \space on \space drives \space gears }"

Gears A,C and E are drives gears while gears B,D, and F are driven gears

So, "\frac{N_A}{N_F}=\frac{T_B*T_D*T_F}{T_A*T_C*T_E}"

"\frac{100}{N_F}=\frac{40*25*24}{60*50*30}"


Angular speed of gear F ,"N_F=375rpm"